OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today?

OpenStudy (anonymous):

First you need to define your unknowns. Can you continue these? Let x be..... Let y be.....

OpenStudy (anonymous):

x = 1 step equations y = 2 step equations x + y = 1150 --> y = 1150 - x 3x - 2y = 1300 now just sub 1150 - x in for y in the second equation to find x can you take it from here ?